OpenStudy (anonymous):

what is the surface area of a sphere with volume 288 pi?

OpenStudy (anonymous):

144pie

OpenStudy (anonymous):

thanks. can yo tell me how you arrived at that answer?

OpenStudy (anonymous):

google

OpenStudy (anonymous):

okay but i have a test tomorrow and I need to know the process involved, not just the answer.

OpenStudy (anonymous):

volume of sphere = 4/3(pi)(r^3) surface area of sphere = 4 (pi) (r^2) hence, 288 = 4/3 r^3 Find what is r then, calculate what is 4 r^2. then u should get the answer of 144 pi helpful?

OpenStudy (anonymous):

yes this helps, thank you